) Please statistically deseribe and determine (statistically test) if the following have equal variances. Also draw a standard ermor interval plot for X and Y (30 pts) 1.4 1.5 2.2 2.4 2.5 2.6 3.5 4.1 5.2 6.3 7.1 7.5
Answer #1

Treatment 1(x)

N1: 6
df1 = N - 1 = 6 - 1 = 5
M1: 2.1
SS1: 1.36
s21 = SS1/(N - 1) = 1.36/(6-1) = 0.27


Treatment 2(y)

N2: 6
df2 = N - 1 = 6 - 1 = 5
M2: 5.62
SS2: 13.17
s22 = SS2/(N - 1) = 13.17/(6-1) = 2.63


T-value Calculation

s2p = ((df1/(df1 + df2)) * s21) + ((df2/(df2 + df2)) * s22) = ((5/10) * 0.27) + ((5/10) * 2.63) = 1.45

s2M1 = s2p/N1 = 1.45/6 = 0.24
s2M2 = s2p/N2 = 1.45/6 = 0.24

t = (M1 - M2)/√(s2M1 + s2M2) = -3.52/√0.48 = -5.05

The t-value is -5.0534. The p-value is .000248. The result is significant at p < .05.

ANSWERED

PLEASE RATE ME POSITIVE

THANKS